OpenStudy (gabylovesyou):

x to the five-fourths power times x to the seven-fourths power, over x to the one-fourth power

OpenStudy (gabylovesyou):

@phi

OpenStudy (gabylovesyou):

\[\frac{ x ^{\frac{ 5 }{ 4}} \times x ^{\frac{ 7 }{ 4 }}}{ x ^{\frac{ 1 }{ 4 }} }\]

OpenStudy (anonymous):

x^12/4 /x^1/4

OpenStudy (phi):

one way to remember what to do is remember this \[ x \cdot x = x^2 \] and x is the same as \( x^1\) so you get this \[ x^1 \cdot x^1= x^2 \] notice the key idea 1+1 = 2

OpenStudy (phi):

add exponents (if you have the same base)

OpenStudy (phi):

and hopefully you know \[ \frac{x\cdot x}{x} = x \] or with exponents \[ \frac{x^2}{x^1} = x^1 \]

OpenStudy (gabylovesyou):

so add 5/4 + 7/4 = 12/4 = 3

OpenStudy (gabylovesyou):

so x^3 ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

exponents that are fractions or negative still work. so yes add

OpenStudy (anonymous):

but x^3 - x^1/4

OpenStudy (phi):

it would be easier to leave it 12/4 because you then subtract off 1/4

OpenStudy (phi):

yes, you get an ugly number \[ x^\frac{11}{4} \]

OpenStudy (gabylovesyou):

Thank you :D